The displacement x of a body varies with time t as x=-2/3 t×t + 16 t +2. In what time of body comes to rest

Grade:11

1 Answers

Vikas TU
7 years ago
velocity should be equal to zero that is dx/dt = 0.
Hence, 
v = dx/dt = -4t/3 + 16 = 0
t/3 = 4
t = 12 s.
After 12 seconds the body will come to rest.